OpenStudy (anonymous):

What is the surface area of this square pyramid? A. 1080 m2 B. 360 m2 C. 504 m2 D. 414 m2

OpenStudy (anonymous):

@taramgrant0543664

OpenStudy (anonymous):

@Shirley27

OpenStudy (xapproachesinfinity):

that pyramid has 4 triangles that are congruent meaning they the same area so 4(the area of one) area of one trianle is 1/2x15x12

OpenStudy (xapproachesinfinity):

add up that with the area of the base, which is a square area of a square = 12^2

OpenStudy (xapproachesinfinity):

that will give the surface area

OpenStudy (xapproachesinfinity):

did yo get it?

OpenStudy (anonymous):

144??

OpenStudy (xapproachesinfinity):

Surface area = 1/2x15x12+12^2

OpenStudy (anonymous):

12 *12 right

OpenStudy (xapproachesinfinity):

don't know i don't have a calculator haha

OpenStudy (xapproachesinfinity):

yes 144 =12^2

OpenStudy (xapproachesinfinity):

you still need to add 1/2 x 15 x12

OpenStudy (xapproachesinfinity):

144+90=?

OpenStudy (anonymous):

got 162

OpenStudy (xapproachesinfinity):

hmm how?

OpenStudy (xapproachesinfinity):

1/2 15 12=90 90+144=234

OpenStudy (anonymous):

\[15\times12\div \frac{ 1 }{ 2}=\]

OpenStudy (anonymous):

+144

OpenStudy (xapproachesinfinity):

no the area of a triangle is 1/2 hxb

OpenStudy (xapproachesinfinity):

\(\frac{1}{2}hb\)

OpenStudy (anonymous):

o srry got 234

OpenStudy (xapproachesinfinity):

ok now you are good!!

OpenStudy (anonymous):

when i did it again

OpenStudy (anonymous):

ok then what

OpenStudy (xapproachesinfinity):

that's all!!

OpenStudy (xapproachesinfinity):

oh hold we forgot to multiple 90 by 4

OpenStudy (anonymous):

brb

OpenStudy (xapproachesinfinity):

because there are four triangle

OpenStudy (xapproachesinfinity):

so 360+144=?

OpenStudy (xapproachesinfinity):

that was blinder from me. i wrote above, i thought i did that already

OpenStudy (xapproachesinfinity):

So surface area =360+144=504 which matches one of your answers

OpenStudy (xapproachesinfinity):

you got it?

OpenStudy (xapproachesinfinity):

if you are confused say so

OpenStudy (anonymous):

no i was doing sumpthin im back now

OpenStudy (xapproachesinfinity):

ok

OpenStudy (anonymous):

thanks can u help with some more

OpenStudy (xapproachesinfinity):

I'm actually leaving to sleep!

OpenStudy (anonymous):

o ok:(((

OpenStudy (xapproachesinfinity):

if you have a quick question n problem

OpenStudy (anonymous):

ok